OpenStudy (anonymous):

Julia has two children who are four years apart in age. Julia is four times older than her youngest child. The sum of the ages of Julia and her children is 76 years. Use the 5‑D Process to find the ages of Julia and each of her children

OpenStudy (caution0):

the kids are a and b julia is j A = B+4 J = 4B A + B + J = 76 (B+4) +B + 4B = 76 6B +4 = 76 6B = 72 B = 12 A = 12+4 = 16 J = 4B = 48